In a class of 500 students, 60 fail in all subjects and the students who pass in exactly 1 subject are twice the students who pass in exactly 2 subjects and thrice the students who pass in all 3 subjects. Find the number of students who fail in atleast 2 subjects.

Grade:12

1 Answers

Rajshree
6 years ago
Let the pass students in one sub. =xN in two=y N three=zSo that X+2x+3x=440X=440/6=220/3Z=3x=3×220/3=220No of students pass at least In two subjects =440-220=220.